OpenStudy (ehsan18):

A postman has to deliver 4 letters to a 4 places, he delivers all of them to the wrong places. What is the number of cases in which it is true?

OpenStudy (anonymous):

suppose A,B,C,D represent the letters mailed in the correct order to the 4 people. How many ways can we send letters so that each one receives the wrong letter B,C,D,A C,D,A,B D,A,B,C B,A,D,C D,C,B,A...

OpenStudy (wolf1728):

In which what is true? You mean in how many ways can he deliver 4 letters to the wrong 4 places? This has to do with a math concept known as "derangements". We could figure this by trial and error. We have four letters which ALL must get delivered to the incorrect address. So let's say letter 1 cannot go to address 1, letter 2 cannot go to address 2 and so on. Basically, we have to determine how many ways we can arrange the numbers 1 2 3 and 4 WITHOUT 1 in place 1, 2 is NOT in place 2, etc. There are NINE such combinations: 2143   2341   2413 3142   3412   3421 4123 4312 4321 Is there a formula? Yes. Number of derangements = n! * (1-1/1! +1/2! -1/3! + 1/4!) Number of derangements = 4! * (1 -1 + (1/2) -(1/6) + (1/24) Number of derangements = 24 * (0 +(1/2) -(1/6) + (1/24) Lowest common denominator of the fractions = 24 Number of derangements = 24 * (0 +(12/24) -(4/24) + (1/24) Number of derangements = 24 * (9/24) Number of derangements = 9 For a calculator and more explanation go here: http://www.1728.org/derange.htm

OpenStudy (wolf1728):

LOL jayzdd gets the medal for an INCORRECT answer ??? !!! The answer is NINE!! LOL

OpenStudy (wolf1728):

Heck, I'LL send a medal to jayzdd too!!!

OpenStudy (anonymous):

I did not finish

OpenStudy (wolf1728):

Well that is okay. I was not really assigning any blame to you - just those medal awarders. (And why are medals awarded BEFORE someone finishes?)

OpenStudy (anonymous):

Using numbers, assuming 1,2,3,4 is the correct order. I get the same 2,1,4,3 2,3,4,1 2,4,1,3 3,1,4,2 3,4,1,2 3,4,2,1 4,1,2,3 4,3,1,2 4,3,2,1

OpenStudy (anonymous):

Nice derangement formula. I did this by making a tree which is tedious

OpenStudy (wolf1728):

Hello to ganeshie8 :-)

OpenStudy (wolf1728):

Or you could have just copied my numbers :-)

OpenStudy (anonymous):